OpenStudy (anonymous):

which value of a solved the equation? a^-2=1/64

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

a - 2 = 1%2F64 ... a = %281%2F64%29+%2B+2 ... a = %281%2F64%29+%2B+2%2864%2F64%29 ... a = %281%2F64%29+%2B+%28128%2F64%29 ... highlight_green%28a+=+129%2F64%29 = 21%2F64

OpenStudy (anonymous):

oops

OpenStudy (the_fizicx99):

1/a^2 = 1/64

OpenStudy (anonymous):

@skullpatrol

OpenStudy (the_fizicx99):

1/8^2 = 1/64

OpenStudy (anonymous):

just go on <SCRIPT LANGUAGE=JavaScript SRC= http://www.algebra.com/cgi-bin/embed-solution.mpl?solution=431220 > </SCRIPT>

OpenStudy (anonymous):

HI :DDDDDD

OpenStudy (anonymous):

what is 2 2/3 - 3/5

OpenStudy (anonymous):

but it is basically 1/a^2 = 1/64

OpenStudy (anonymous):

2.06666666667

OpenStudy (the_fizicx99):